OpenStudy (anonymous):

Find the first six terms of the sequence. (1 point) a1 = 6, an = 2 • an-1 (A) 12, 24, 48, 96, 192, 384 (B) 6, 12, 14, 16, 18, 20 (C) 0, 2, 12, 14, 16, 18 (D) 6, 12, 24, 48, 96, 192

OpenStudy (anonymous):

You're given the first term, seems you just have to plug in other terms into your formula you're given.

OpenStudy (anonymous):

How can I set that up? Take me through the first one as an example @iambatman

OpenStudy (anonymous):

First, tell me what your first term is, then we can eliminate two of the choices.

OpenStudy (anonymous):

That's literally the entire question word for word, there isn't anything else @iambatman

OpenStudy (anonymous):

I'm asking you for the first term.

OpenStudy (anonymous):

What is it, it's given.

OpenStudy (anonymous):

6?

OpenStudy (anonymous):

Yes, that's what a1 means, and I'm assuming your problem is actually \[a_n=2 \times a_n-1\]

OpenStudy (anonymous):

yes @iambatman

OpenStudy (anonymous):

Ok so plug in the value 2, since we're looking for a_2

OpenStudy (anonymous):

That will give you a_2 = 2*a1 => a_2 = 2*6 = 12

OpenStudy (anonymous):

Do the same with the rest of them

OpenStudy (anonymous):

oh okay thank you @iambatman